Let Ω be the region above y+z= 2, below z= 4, and between x= 0 and x= 4−y^2 a triple integral for the volume of the region. Set up an integral to represent the volume of Ω in all six orders and then fully solve ONE of them.
1
Expert's answer
2019-10-07T10:04:46-0400

There are 6 possible ordersof integration


"dzdxdy,dzdydx,dydzdx,dydxdz,dxdydz,dxdzdy"

"V=\iiint_\Omega dV"

"-2\leq y\leq 2, \ 0\leq x\leq 4-y^2,\ 2-y\leq z\leq 4"

"V=\displaystyle\int_{-2}^2\displaystyle\int_{0}^{4-y^2}\displaystyle\int_{2-y}^4dzdxdy="

"=\displaystyle\int_{-2}^2\displaystyle\int_{0}^{4-y^2}[z]\begin{matrix}
 4 \\
 2-y
\end{matrix}dxdy=\displaystyle\int_{-2}^2\displaystyle\int_{0}^{4-y^2}(2+y)dxdy="

"=\displaystyle\int_{-2}^2[2x+xy]\begin{matrix}
 4-y^2 \\
 0
\end{matrix}dy=\displaystyle\int_{-2}^2\big(2(4-y^2)+y(4-y^2)-0\big)dy="

"=\bigg[ 8y-{2y^3 \over 3}+2y^2-{y^4 \over 4}\bigg]\begin{matrix}
 2 \\
 -2
\end{matrix}="

"=16-{16 \over 3}+8-4+16-{16 \over 3}-8+4={64 \over 3}(units^3)"

"0\leq x\leq 4, \ -\sqrt{4-x}\leq y\leq \sqrt{4-x},\ 2-y\leq z\leq 4""V=\displaystyle\int_{0}^4\displaystyle\int_{-\sqrt{4-x}}^{\sqrt{4-x}}\displaystyle\int_{2-y}^4dzdydx"



"-2\leq y\leq 2, \ 2-y\leq z\leq 4,\ 0\leq x\leq 4-y^2""V=\displaystyle\int_{-2}^2\displaystyle\int_{2-y}^4\displaystyle\int_{0}^{4-y^2}dxdzdy"



"0\leq z\leq 4, \ -2\leq y\leq 2-z,\ 0\leq x\leq 4-y^2""V=\displaystyle\int_{0}^4\displaystyle\int_{-2}^{2-z}\displaystyle\int_{0}^{4-y^2}dxdydz"



"0\leq z\leq 4, \ 0\leq x\leq 4-(2-z)^2,\ -2\leq y\leq 2-z""V=\displaystyle\int_{0}^4\displaystyle\int_{0}^{4-(2-z)^2}\displaystyle\int_{-2}^{2-z}dydxdz"



"0\leq x\leq 4, \ 2-\sqrt{4-x}\leq z\leq 2+\sqrt{4-x},\ -2\leq y\leq 2-z""V=\displaystyle\int_{0}^4\displaystyle\int_{2-\sqrt{4-x}}^{2+\sqrt{4-x}}\displaystyle\int_{-2}^{2-z}dydzdx"